OpenStudy (anonymous):

NEEED HELP PLS x^2+7x+1=-9 factor left hand side of equation into two linear equations

OpenStudy (anonymous):

ill do anything for some help please

OpenStudy (anonymous):

YES I WILL GIVE YOU THE TICKETS TO LAS VEGAS AND A DELIVERY OF YOUR BRAND NEW BMW filled with 10,000 all in one dollar bills d:

OpenStudy (shadowlegendx):

Cant forget the cheese burger...

OpenStudy (anonymous):

oh and a cheeseburger , im serious too!!!!!

OpenStudy (whpalmer4):

\[x^2+7x+1=-9\]\[x^2+7x+1+9=-9+9\]\[x^2+7x+10=0\] Now pick two factors of 10 that add to 7. Call them \(a,b\). Your factoring is\[x^2+7x+10 = (x+a)(x+b)\]

OpenStudy (anonymous):

ugh call me a a cab and you got yourself a deal lol.. and whpalmer would it be 2 and 5?

OpenStudy (whpalmer4):

Let's check! \[(x+2)(x+5) = x(x+5) + 2(x+5) = x*x + 5*x + 2*x + 2*5\]\[ = x^2 + 7x + 10\checkmark \]

OpenStudy (anonymous):

the biggest problem i have is i do not know how to write it

OpenStudy (anonymous):

do i write it like (x+5) (x+2) -_-

OpenStudy (whpalmer4):

Yes, that's your two linear equations. No exponents other than 1 on the \(x\)'s, so they are linear.

OpenStudy (whpalmer4):

something like \((x^2+1)\) would not be linear.